three years ago Tom was twice as old as Jean, and in two years the sum of their ages will be 28 years. Find the present ages?
Answer by ankor@dixie-net.com(22740)

let t = Tom's present age
let j = Jean's
:
Write an equation for each statement,simplify as much as possible
:
three years ago Tom was twice as old as Jean,
t-3 = 2(j-3)
t - 3 = 2j - 6
t = 2j - 6 + 3
t = 2j - 3
and in two years the sum of their ages will be 28 years.
(t+2) + (j+2) = 28
t + j + 4 = 28
t + j = 28 - 4
t + j = 24
Find the present ages?
replace t with (2j-3) from the 1st statement
2j - 3 + j = 24
2j + j = 24 + 3
3j = 27
j = 27/3
j = 9 yrs is Jean's age
Find Tom's age
t = 2(9) - 3
t = 15 yrs is Tom's age
:
:
Check this in the first statement
" three years ago Tom was twice as old as Jean, "
15 - 3 = 2(9 - 3)
12 = 2(6)